OpenStudy (yacoub1993):

Find the slope of the line through the pair of points. A(2, –3), P(2, 9) A. 0 B. 2/3 C. - 1/3 -1/3 D. undefined

OpenStudy (anonymous):

use slope formula : for two points did u knw that may i tell u ?

hero (hero):

y = mx + b -3 = 2m + b 9 = 2m + b b = -3 - 2m b = 9 - 2m b = b -3 - 2m = 9 - 2m 2m - 2m = 9 - 3 0 = 6 False

OpenStudy (yacoub1993):

thnxs guys

hero (hero):

Bro do you know which is the correct answer choice?

OpenStudy (yacoub1993):

D IS THE CORRECT ANSWER CHOICE sry for late answering

hero (hero):

Good job

OpenStudy (yacoub1993):

now can answer the other question